Paul and his brothers Ted and Al worked out one day that their father was 3 times as old as Ted. Also that, in three years time, he would be three times as old as Paul would be, and that in 6 years time he would be twice as old as Al would be. Al was two years older than Ted.
How old is their father?

Answer by ankor@dixie-net.com(22740)

Write an equation for each statement:
:
Paul and his brothers Ted and Al worked out one day that their father was 3 times as old as Ted.
f + 3t
:
Also that, in three years time, he would be three times as old as Paul would be,
f + 3 = 3(p+3)
f + 3 = 3p + 9
f = 3p + 9 - 3
f = 3p + 6
:
and that in 6 years time he would be twice as old as Al would be.
f + 6 = 2(a+6)
f + 6 = 2a + 12
f = 2a + 12 - 6
f = 2a + 6
:
Al was two years older than Ted.
a = t + 2
In the equation f = 2a + 6 replace f with 3t
3t = 2a + 6
Replace a with (t+2),find t
3t = 2(t+2) + 6
3t = 2t + 4 + 6
3t - 2t = 10
t = 10 yrs is Teds present age
then
3(10) = 30 yrs is father's age, which all they wish to know
: